# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v2

@@ Abstract.
Online Communities, Reddit, Rules, AIGC, Generative AI, Governance, AI-Generated Content, Moderation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
The emergence of mainstream generative AI tools capable of producing compelling text, images, and videos—collectively referred to as AI-Generated Content (AIGC)—poses a unique challenge for online platforms that are dependent on User-Generated Content. Many such platforms have enacted top-down changes to respond to this new technology, with some deciding to ban it (Makyen, 2023), while others are encouraging its use or even incorporating generative AI features into their user interfaces (Lunden, 2023; Mehta, 2023). Other platforms have declined to make top-down changes and have instead left it up to their users to come to their own positions on this technology. One such platform is Reddit, a social sharing and news aggregation site, where individual communities, known as subreddits, are free to enact their own rules and policies (Fiesler et al., 2018; Proferes et al., 2021; Singer et al., 2014; Gilbert, 2013). This decentralized moderation approach allows communities to craft policies towards generative AI that are contextually-relevant: the policies align with how communities see the technology being used, and whether or not that use aligns with their values. Decentralized moderation also provides a rich source of data through which HCI researchers interested in helping online communities navigate sudden technological change can better understand the range of attitudes and stances that exist across communities who vary in size, purpose, and values.
@@
Findings from past HCI studies have emphasized the important role that explicit and consistent community rules play in helping online communities navigate moments of intense change (Kiene et al., 2016). These rules interact with other forms of governance (Lessig, [n. d.]), such as norms (Chandrasekharan et al., 2018) and technical affordances (Jhaver et al., 2019), to encourage certain behaviors and discourage others (Kiesler et al., 2012).
@@
Of course, generative AI and AIGC may impact the rules (Makyen, 2023), dynamics (Hua et al., 2024; Fiesler, 2024), and governance (Cao et al., 2023) strategies of online communities in complex ways. While some recent research has explored how generative AI might help online communities deal with problematic content (Agarwal et al., 2024; Kumar et al., 2024), other work has documented the ability of AIGC to pollute these same communities with deceptive information (Park et al., 2024). Recent work by Lloyd et al. (Lloyd et al., 2023) used interviews with Reddit moderators to explore the particular challenges of governing AI use in online communities and found that communities are enacting explicit rules about AIGC in order to clarify norms. Lloyd et al.’s work focuses on the perspectives and experiences of their interviewees and raises the question of how their findings generalize across the wide variety of communities on Reddit.
@@
This paper addresses that question through a large-scale quantitative analysis of community rules on Reddit. Our approach allows us to paint a fuller picture for the HCI community by complementing Lloyd et al.’s qualitative findings with an analysis of platform-wide patterns and trends, while simultaneously providing deeper insight into the specific types of AI rules that communities are creating. We build on techniques developed by Fiesler et al. (2018) to study the ecosystem of community rules on Reddit and adapt their methodology to specifically investigate the emergence of community rules governing AI use. We use Fiesler et al.’s community topic and rule taxonomies so that we can compare our analysis of AI rules to their analysis of general Reddit rules. We address the following research questions:
@@
To answer these questions we crawled over $300,000$ public subreddits on two occasions, in July 2023 and November 2024, building a dataset of $N=99,969$ English-language subreddits with published rules. Analyzing a longitudinal sample of subreddits that were present during both our crawls, our inquiry into RQ1 finds that while only $1.2\%$ of subreddits now have rules about AI, the number of subreddits with AI rules has more than doubled since July 2023. Larger subreddits are much more likely to have AI rules: for example, of the top $1\%$ largest subreddits, $17\%$ had AI rules in our most recent crawl. To explore RQ2, we perform subreddit-level analysis on all subreddits in our 2024 sample with rules. We apply a “Topic” label to these subreddits using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and “Community Archetype” labels using a taxonomy from Prinster et al. (Prinster et al., 2024). We use quantitative methods to explore the relationship between these labels and the presence of AI rules and find these rules to be most common in communities with Art and Celebrity topics, as well as those that fit the Content Generation Community Archetype. To explore RQ3, we perform rule-level analysis on the rules in our sample that we identified as governing AI use. We apply an existing rule taxonomy from Fiesler et al. (Fiesler et al., 2018) and additionally produce our own set of emergent labels to describe meaningful dimensions of variation in both the requirements imposed by these rules and the language through which they frame AI usage. We find rules about AI images to be more common than rules about other types of AIGC and see quality and authenticity as the most frequently expressed concerns.
We compare these findings to the results of related work and discuss implications for platform designers. In particular, thoughtful design choices may help address the community concerns about quality and authenticity that our findings identify. Our findings also show, though, that communities of different types take varied approaches towards regulating AI use, and suggest that design solutions meant to help communities address these challenges must be sensitive to community context.
As an additional contribution, we make our community rules data and labels public. While we limit our current analysis to questions about AI rules, our data can be used to answer a wide range of questions about content moderation in self-moderating online communities. We hope that this will enable large-scale analysis by future HCI researchers interested in encouraging online community self-determination.
@@ 2. Related Work
Our study is motivated by two bodies of work from the HCI literature. The first set of related work includes empirical research into online communities’ rules, norms, and values. This work demonstrates that online communities hold a wide range of values and emphasizes that rules and norms are always context-dependent. The second body of work concerns current and potential future impacts of generative AI on online communities. This research shows that communities are already grappling with the changes brought by generative AI and are adapting their policies and practices in response.
@@ 2.1. Online Communities’ Rules, Norms, and Values
Online communities play an important role in social life (Preece, 1998), and occupy a complicated position as a site of nuanced interactions between people, technology, and policy (Jackson et al., 2014). This importance and complexity have made online communities a frequent topic of study in the HCI literature (Preece, 2000; Kraut and Resnick, 2012). HCI scholars have explored the variety of purposes that these communities can serve and how platform design choices may influence their success. For example, Preece introduced the idea of Empathic Communities (Preece, 1999), in which people come together to share and receive support, and suggested platform design guidelines to encourage sociality (Preece, 2001). In addition to encouraging pro-social behavior, a frequent topic of study has been how to discourage anti-social behavior (Kiesler et al., 2012; Lessig, [n. d.]). Much of this work has focused on the benefits (Grimmelmann, 2015) and challenges (Koshy et al., 2023) of effective and fair content moderation. The HCI community has been interested in top-down moderation (Gillespie, 2018), in which platforms set and enforce rules, but even more interested in bottom-up, or self-moderation, in which communities do (Seering, 2020). A major finding of this research is the importance of context sensitivity in moderation decisions, which self-moderation is often more attuned to. Taken together, this research suggests that studying self-moderation decisions may provide insight for design solutions that can promote community self-determination at a time of sudden technological change.
@@
Reddit is a fruitful platform for a study of community self-moderation because it is home to online communities of various size and purpose. This heterogeneity, combined with its largely public and text-based nature, have made Reddit a frequent topic of study in online communities research (Proferes et al., 2021). Recent work has explored the values (Weld et al., 2024), norms (Chandrasekharan et al., 2018), and rules (Fiesler et al., 2018) of different communities across the platform, both via direct methods like surveys, and indirect methods such as analyzing digital trace data from these communities. Reddit’s implementation of community self-moderation allows communities to write and make public their own community-enforced rules, which Fiesler et al. collect and characterize in a study that is the main methodological inspiration for our work (Fiesler et al., 2018). In their study, Fiesler et al. find subreddit size to be the main predictor of having rules, but note that even among subreddits of similar sizes rules are context-dependent, not uniform, and often uniquely tailored to the topic of the subreddit (e.g., art subreddits have more rules about copyright). The authors also highlight that setting norms and developing rules is fundamentally a social process. Overall, Fiesler et al. show the importance of community rules and the feasibility of using them as a data source to study attitudes and stances across a platform. Given the potential impact of AI on online communities, we use community rules as a starting point to study context-specific attitudes towards AIGC as they emerge across Reddit.
@@ 2.2. Generative AI’s Impact on Online Communities
A growing body of research in HCI, Social Computing, and adjacent behavioral science fields has documented ways in which generative AI may impact online communities through increased deception, decreased trust, and new forms of harassment. Studies have demonstrated that AI can effectively deceive humans (Park et al., 2024) and have raised concerns that generated text could be weaponized for deception by either humans (Goldstein et al., 2023) or automated social media accounts (bots) (Menczer et al., 2023). Additionally, just the suspicion that other online community members are using AI may be enough to impact dynamics in online communities. Several studies in AI-Mediated Communication (AIMC) (Hancock et al., 2020) provide evidence that the perception that another online community member is using AI causes members to regard each other more negatively (Rae, 2024) and trust each other less (Jakesch et al., 2019). Finally, HCI research has long documented how online communities can be sites of harassment and abuse (Jiang et al., 2019; Park et al., 2022; Cheng et al., 2017; Han et al., 2023). The impact of AI-powered harassment on online communities, such as malicious synthetic, or deepfake, media (Farid, 2022) is much less studied (Umbach et al., 2024), but there is concern that this extremely harmful phenomenon may grow in prevalence alongside the popularity of generative AI tools (Rini and Cohen, 2022).
@@
Many additional studies have started to document the impacts of AIGC on online communities “in the wild”. Recent work has identified undisclosed AI use in several settings, such as LLM-powered social bots on Twitter (Yang and Menczer, 2024), AIGC in fake online reviews (Oak and Shafiq, 2024), and spam content used for scams (DiResta and Goldstein, 2024). In addition to studies of deceptive use, other recent work has begun to explore how the disclosed use of AIGC may affect community dynamics, such as in art-sharing communities (Wei and Tyson, 2024). Generative AI can additionally impact online communities by drawing traffic away from communities and decreasing overall engagement: causal inference studies of Q&A sites in the Stack Exchange network, where users gather to ask and answer questions about programming and related topics, have demonstrated that the launch of ChatGPT caused an overall decline in website visits and question volume (Burtch et al., 2024). The authors of this study contrast this decline with a measured null-effect on similar Reddit communities, which they suggest may avoid a decline in traffic because of their emphasis on socialization, rather than pure information exchange. Finally, recent work by Lloyd et al. used interviews to qualitatively explore the attitudes and experiences of Reddit moderators dealing with AIGC (Lloyd et al., 2023). This study interviewed moderators from subreddits of different sizes and topics and surfaces a variety of community objections to AIGC that range from practical to ideological. The work poses open questions about how prevalent these concerns are across the platform and suggests that larger trends could be explored quantitatively in future work. Our study fills that gap through a quantitative analysis that helps identify the most common rules and the most impacted communities, so that the HCI community can begin the task of addressing these issues with context-specific solutions.
@@ 3. Methods
In order to perform a large-scale analysis of community rules governing the use of AI, we collected the metadata of public, English-language subreddits. Building on a large-scale crawl in 2024, and an earlier one performed in 2023, we created five datasets of subreddits and their associated rules. We processed the more recent crawl to create four datasets: the Broad Subreddit Set ($N=307,543$) consisting of the metadata of subreddits with and without rules; the Rules Subreddit Set ($N=99,969$) of only subreddits with stated, published rules; the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing the use of AI; and the AI Rules Set ($N=4,458$) of individual rules governing AI use. The process of constructing these datasets is summarized in Figure 1, and detailed below. We supplement the raw crawl data with additional information about subreddit and rule types through several rounds of labeling, using both human and LLM coders. We use the AI Rules Set and emergent coding to construct a novel taxonomy of AI Rule Types. In addition, we created a longitudinal dataset of subreddits that were present in both the 2023 and 2024 crawls. We use this Longitudinal Subreddit Set ($N=227,737$) to compare the prevalence of AI rules at different points in time for a consistent sample of subreddits. We will make our datasets and labels public upon completion of peer-review.
Figure 1. The process for collecting subreddit metadata and building the datasets used in our analysis: the Broad Subreddit Set of English-language subreddits, the Rules Subreddit Set of subreddits with non-empty rules, the AI Rules Subreddit Set of subreddits with rules governing the use of AI, and the AI Rules Set of individual rules governing AI use.
@@
A funnel diagram with five steps showing the filtering of subreddits based on various criteria: 1. Step 1: 331,938 subreddits, representing 100% of all subreddits seen during the 2024 crawl. 2. Step 2: 307,543 subreddits (92.7%), filtered to subreddits older than 1 month old where English is the primary language. 3. Step 3: 99,969 subreddits (30.1%), filtered to subreddits with non-empty ”rules”. 4. Step 4: 4,251 subreddits (23.8%), further filtered to subreddits with rules governing AI use. The final output is 4,458 individual AI rules identified in these subreddits.

### v3

@@ Abstract.
Online Communities, Reddit, Rules, Generative AI, Governance, AI-Generated Content, Moderation
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713292isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
The emergence of mainstream generative AI tools capable of producing compelling text, images, and videos—collectively referred to as AI-generated content (AIGC)—poses a unique challenge for online platforms that are dependent on user-generated content. Many such platforms have enacted top-down changes to respond to this new technology, with some banning it (Makyen, 2023), while others are encouraging its use or incorporating generative AI features into their user interfaces (Lunden, 2023; Mehta, 2023). Other platforms have declined to make top-down changes and have instead left it up to users to come to their own positions on this technology. One such platform is Reddit, a social sharing and news aggregation site, where individual communities, known as subreddits, are free to enact their own rules and policies (Fiesler et al., 2018; Proferes et al., 2021; Singer et al., 2014; Gilbert, 2013). This decentralized moderation approach allows communities to craft generative AI policies that are contextually-relevant: the policies are based on how communities observe the technology being used, and whether or not that use aligns with their values. Decentralized moderation also provides a rich source of data through which HCI researchers interested in helping online communities navigate sudden technological change can better understand the range of attitudes and stances that exist across communities who vary in size, purpose, and values.
@@
Findings from past HCI studies have emphasized the important role that explicit and consistent community rules play in helping online communities navigate moments of intense change (Kiene et al., 2016). These rules interact with other forms of governance (Lessig, 1999), such as norms (Chandrasekharan et al., 2018) and technical affordances (Jhaver et al., 2019), to encourage certain behaviors and discourage others (Kiesler et al., 2012).
@@
Of course, generative AI and AIGC may impact the rules (Makyen, 2023), dynamics (Hua et al., 2024; Fiesler, 2024), and governance (Cao et al., 2024) strategies of online communities in complex ways. While some recent research has explored how generative AI might help online communities deal with problematic content (Agarwal et al., 2024; Kumar et al., 2024), other work has documented the ability of AIGC to pollute these same communities with deceptive information (Park et al., 2024). In our own earlier work (Lloyd et al., 2024), we used interviews with Reddit moderators to explore the particular challenges of governing AI use in online communities and found that communities are enacting explicit rules about AIGC in order to clarify norms. This earlier study focused on the perspectives and experiences of the interviewees and raised the question of how the findings generalize across the wide variety of communities on Reddit.
@@
This paper expands our understanding of Reddit communities’ responses to AIGC through a large-scale quantitative analysis of community rules. Our approach allows us to paint a fuller picture for the HCI community by complementing the earlier qualitative findings with an analysis of platform-wide patterns and trends, while simultaneously providing deeper insight into the specific types of AI rules that communities are creating. We build on techniques developed by Fiesler et al. (2018) to study the ecosystem of community rules on Reddit and adapt their methodology to specifically investigate the emergence of community rules governing AI use. We use Fiesler et al.’s taxonomies for community topics and rules so that we can compare our analysis of AI rules to their analysis of general Reddit rules. We address the following research questions:
@@
To answer these questions we crawled over $300,000$ public subreddits on two occasions, in July 2023 and November 2024, building a dataset of $N=99,969$ English-language subreddits with published rules. Analyzing a longitudinal sample of subreddits that were present during both our crawls, our inquiry into RQ1 finds that while in November 2024 only $1.2\%$ of subreddits had rules about AI, the number of subreddits with AI rules more than doubled since July 2023. Larger subreddits are much more likely to have AI rules: for example, of the top $1\%$ largest subreddits in our 2024 sample, $17\%$ had AI rules. To explore RQ2, we perform subreddit-level analysis on all subreddits with rules in our 2024 sample. We apply a “Topic” label to these subreddits using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and “Community Archetype” labels using a taxonomy from Prinster et al. (Prinster et al., 2024). We use quantitative methods to explore the relationship between these labels and the presence of AI rules and find these rules to be most common in communities with Art and Celebrity topics, as well as communities that fit the Content Generation Community Archetype. To explore RQ3, we perform rule-level analysis on the rules in our sample that we identified as governing AI use. We apply an existing rule taxonomy from Fiesler et al. (Fiesler et al., 2018) as well as a novel AI Rules Taxonomy to describe meaningful dimensions of variation in both the requirements imposed by these rules and the language through which they frame AI usage. We find that rules about AI images are more common than rules about other types of AIGC, and see quality and authenticity as the most frequently mentioned concerns.
Our findings offer implications for designers of online community platforms. In particular, community concerns about quality and authenticity may be addressed using thoughtful design choices that encourage effortful communication (Kelly et al., 2017; Zhang et al., 2022). Our findings also show, though, that communities of different types take varied approaches towards regulating AI use, and suggest that design solutions meant to help communities address these challenges must be sensitive to community context.
As an additional contribution, we make our community rules data and labels public [^1] https://github.com/sTechLab/AIRules. While we limit our current analysis to questions about AI rules, our data can be used to answer a wide range of questions about governance in self-moderating online communities. We hope that our data will enable large-scale analysis by future HCI researchers interested in encouraging online community self-determination.
@@ 2. Related Work
Our study is motivated by two bodies of work from the HCI literature. The first includes empirical research into online communities’ rules, norms, and values. This work demonstrates that online communities hold a wide range of values and emphasizes that rules and norms are context-dependent. The second body of work concerns current and potential future impacts of generative AI on online communities. This research shows that communities are already grappling with the changes brought by generative AI and are adapting their policies and practices in response.
@@ 2.1. Online Communities’ Rules, Norms, and Values
Online communities play an important role in social life (Preece, 1998), and occupy a complicated position as a site of nuanced interactions between people, technology, and policy (Jackson et al., 2014). This importance and complexity have made online communities a frequent topic of study in the HCI literature (Preece, 2000; Kraut and Resnick, 2012). HCI scholars have explored the variety of purposes that these communities can serve and how platform design choices may influence their success. For example, Preece introduced the idea of Empathic Communities (Preece, 1999), in which people come together to share and receive support, and suggested platform design guidelines to encourage sociality (Preece, 2001). In addition to encouraging pro-social behavior, a frequent topic of study has been how to discourage anti-social behavior (Kiesler et al., 2012; Lessig, 1999). Much of this work has focused on the benefits (Grimmelmann, 2015) and challenges (Koshy et al., 2023) of effective and fair content moderation. The HCI community has been interested in top-down moderation (Gillespie, 2018), in which platforms set and enforce rules, as well as bottom-up, or self-moderation, in which communities perform these tasks (Seering, 2020). A major result of this body of research is a recognition of the importance of context sensitivity in moderation decisions, which self-moderation is often more attuned to. Taken together, this research suggests that studying self-moderation decisions may provide insight for design solutions that can promote community self-determination in times of sudden technological change.
@@
Reddit is a useful site for a study of community self-moderation because it is home to online communities of various size and purpose. This heterogeneity, combined with its largely public and text-based nature, have made Reddit a frequent topic of study in online communities research (Proferes et al., 2021). Recent work has used surveys and digital trace data to explore the values (Weld et al., 2024), norms (Chandrasekharan et al., 2018), and rules (Fiesler et al., 2018) of different communities across the platform. Reddit’s implementation of community self-moderation allows communities to write and make public their own community-enforced rules, which Fiesler et al. collect and characterize in a study that is the main methodological inspiration for our work (Fiesler et al., 2018). In their study, Fiesler et al. find subreddit size to be the main predictor of having rules, but note that even among subreddits of similar sizes rules are context-dependent, not uniform, and often uniquely tailored to the topic of the subreddit (e.g., art subreddits have more rules about copyright). The authors also highlight that setting norms and developing rules is fundamentally a social process. Overall, Fiesler et al. show the importance of community rules and the feasibility of using them as a data source to study attitudes and stances across a platform. Given the potential impact of AI on online communities, we use community rules as a starting point to study context-specific attitudes towards AIGC as they emerge across Reddit.
@@ 2.2. Generative AI’s Impact on Online Communities
A growing body of research in HCI, Social Computing, and adjacent behavioral science fields has documented ways in which generative AI may impact online communities by increasing deception, decreasing trust, and enabling new forms of harassment. Studies have demonstrated that AI can effectively deceive humans (Park et al., 2024) and have raised concerns that generated text could be weaponized for deception by either humans (Goldstein et al., 2023) or automated social media accounts (bots) (Menczer et al., 2023). Additionally, just the suspicion that other online community members are using AI may be enough to impact dynamics in online communities. Several studies in AI-Mediated Communication (AIMC) (Hancock et al., 2020) provide evidence that the perception that another online community member is using AI causes members to regard each other more negatively (Rae, 2024) and trust each other less (Jakesch et al., 2019). Finally, HCI research has long documented how online communities can be sites of harassment and abuse (Jiang et al., 2019; Park et al., 2022; Cheng et al., 2017; Han et al., 2023). The impact of AI-powered harassment on online communities, such as malicious synthetic, or deepfake, media (Farid, 2022) is much less studied (Umbach et al., 2024), but there is concern that this extremely harmful phenomenon may grow in prevalence alongside the popularity of generative AI tools (Rini and Cohen, 2022).
@@
Many additional studies have started to document the impacts of AIGC on online communities “in the wild”. Recent work has identified undisclosed AI use in several settings, such as LLM-powered social bots on Twitter (Yang and Menczer, 2024), AIGC in fake online reviews (Oak and Shafiq, 2024), and in spam content used for scams (DiResta and Goldstein, 2024). In addition to studies of deceptive use, other recent work has begun to explore how the disclosed use of AIGC may affect community dynamics, such as in art-sharing communities (Wei and Tyson, 2024; Matatov et al., 2024). Generative AI can additionally impact online communities by drawing traffic away and decreasing overall engagement: studies of Q&A sites in the Stack Exchange network, where users gather to ask and answer questions about programming and related topics, have demonstrated that the launch of ChatGPT caused an overall decline in website visits and question volume (Burtch et al., 2024). The authors of this study contrast this decline with a measured null-effect on similar Reddit communities, which they suggest may avoid a decline in traffic because of their emphasis on socialization, rather than pure information exchange. Finally, our previous work used interviews to qualitatively explore the attitudes and experiences of Reddit moderators dealing with AIGC (Lloyd et al., 2024). In that study, we interviewed moderators from subreddits of different sizes and topics. This earlier work surfaces a variety of community objections to AIGC that range from practical to ideological. The findings of this qualitative study pose open questions about how prevalent these concerns are across the platform and suggest that these trends could be explored quantitatively in future work. This current study fills that gap through a quantitative analysis that helps identify the most common rules and the most impacted communities, so that the HCI community can begin the task of addressing these issues with context-specific solutions.
@@ 3. Methods
In order to perform a large-scale analysis of community rules governing the use of AI, we collected the metadata of public, English-language subreddits. Building on a large-scale crawl in November 2024, and an earlier one performed in July 2023, we created five datasets of subreddits and their associated rules. We processed the more recent crawl to create four datasets: the Broad Subreddit Set ($N=307,543$) consisting of the metadata of subreddits with and without rules; the Rules Subreddit Set ($N=99,969$) of only subreddits with stated, published rules; the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing the use of AI; and the AI Rules Set ($N=4,458$) of individual rules governing AI use. The process of constructing these datasets is summarized in Figure 1, and detailed below. We supplement the raw crawl data with additional information about subreddit and rule types through several rounds of labeling, using both human and LLM coders. We use emergent coding to construct a novel taxonomy of AI Rule Types from the AI Rules Set. In addition, we created a longitudinal dataset of subreddits that were present in both the 2023 and 2024 crawls. We use this Longitudinal Subreddit Set ($N=227,737$) to compare the prevalence of AI rules at different points in time for a consistent sample of subreddits. Our datasets and labels are publicly available at https://github.com/sTechLab/AIRules.
Figure 1. The process for collecting subreddit metadata and building the datasets used in our analysis.
@@
A funnel diagram with four steps showing the filtering of subreddits based on various criteria: 1. Step 1: 331,938 subreddits, representing 100% of all subreddits seen during the 2024 crawl. 2. Step 2: 307,543 subreddits (92.7%), filtered to subreddits older than 1 month old where English is the primary language. 3. Step 3: 99,969 subreddits (30.1%), filtered to subreddits with non-empty ”rules”. 4. Step 4: 4,251 subreddits (1.28%), further filtered to subreddits with rules governing AI use. The final output is 4,458 individual AI rules identified in these subreddits.